OpenStudy (anonymous):

What is the x-coordinate of the vertex of the parabola whose equation is y = 3x^2 + 9x? a) -3 b) -1 1/2 c) -2/3

OpenStudy (anonymous):

I got -3/2.

OpenStudy (anonymous):

\(-\frac{b}{2a}\) which in your case is \(-\frac{9}{2\times 3}\)

OpenStudy (anonymous):

so yeah, \(-\frac{3}{2}\)